OpenStudy (anonymous):

One airplane is approaching an airport from the north at 181km/hr. Asecond airplane approaches from the east at 278km/hr. Find the rate at which the distance between the planes change when the southbound plane is 30km away from the airport and the westbound plane is 15km from the airport.

OpenStudy (anonymous):

\[z ^{2} =\sqrt{x ^{2} + y ^{2} } = \sqrt{15 ^{2} + 30 ^{2} } = 15\sqrt{5}\] \[z' = (xx' + yy')/z = (15* 278 + 30 * 181)/ 15\sqrt{5} = 286 km/h\]